# Combinatorics of Essential Sets for Positroids

### v2

@@ 5 Realization spaces of positroids
We now turn our attention to the realization space of a positroid in the totally non-negative Grassmannian and the Grassmannian.
Recall that, given a field $\mathbb{K}$, the Grassmannian $\operatorname{Gr}(k,n)$ is the space of all $k$-dimensional linear subspaces of $\mathbb{K}^{n}$. Any point $V$ in $\operatorname{Gr}(k,n)$ can be represented by a $k\times n$ matrix with entries in $\mathbb{K}$. Let $M=(m_{ij})$ be a $k\times n$ matrix of indeterminates. For a subset $I=\{i_{1},\ldots,i_{k}\}\in\binom{[n]}{k}$, let $M_{I}$ denote the $k\times k$ submatrix of $M$ with the column indices $i_{1},\ldots,i_{k}$. The Plücker coordinates of $V$ are $p_{I}(V)=\text{det}(M_{I})$ for $I\in\binom{[n]}{k}$ that do not depend on the choice of matrix $M$ (up to simultaneous rescaling by a non-zero constant) and determine the Plücker embedding of $\operatorname{Gr}(k,n)$ into $\mathbb{P}^{\binom{n}{k}-1}$. Moreover, any point $V\in\operatorname{Gr}(k,n)$ can be represented as $\text{span}\{v_{1},\ldots,v_{k}\}$ for some $\mathbb{K}$-vector space basis $\{v_{1},\dots,v_{k}\}$.
@@ 5.1 Positroid cells
Let $\mathbb{K}=\mathbb{R}$. The totally non-negative Grassmannian $\operatorname{Gr}_{\geq 0}(k,n)$ is the subset of $\operatorname{Gr}(k,n)$ consisting of points whose Plücker coordinates have all the same sign. Fix a positroid $\mathcal{P}$ on $[n]$ of rank $k$, defined via its basis $\mathcal{B}$. Then the positroid cell of $\mathcal{P}$ is the set:
$\mathcal{S}_{\mathcal{P}}=\{M\in\operatorname{Gr}_{\geq 0}(k,n)\mid M_{I}=0\text{ if and only if }I\neq B,\text{ for all }B\in\mathcal{B}\}.$
@@
Let $\mathcal{E}^{r}$ be the set of ranked essential sets of a rank $k$ positroid $\mathcal{P}$. Then
$\operatorname{codim}(\mathcal{S}_{\mathcal{P}})=\sum_{(r,[i,j])\in\mathcal{E}^{r}}e_{[i,j]}(k-r)=\sum_{(r,[i,j])\in\mathcal{C}^{r}}e_{[i,j]}(k-r).$
@@
Let $\pi$ be the bounded affine permutation corresponding to the positroid $\mathcal{P}$. By Theorem 5.1, we need to prove that $\ell(\pi)=\sum_{(r,[i,j])\in\mathcal{E}^{r}}e_{[i,j]}(k-r)$. Consider the dotted array of $\pi$ and note that the length of $\pi$ can be computed by looking at the array as:
@@
Note that, if the dot in row $i$ lies in $D(\pi)\cap T_{(j,\pi(j)-j+1)}$, then $(j,\pi(j)-j+1)$ defines a bounded section of the diagram of $\pi$, which is non-empty since it contains the dot $(i,\pi(i)-i+1)$. Hence there is a corner, i.e., an essential set inside $P_{(j,\pi(j)-j+1)}$. Let $(r,E)\in\mathcal{E}^{r}$ be such an essential set. The number of dots in the same connected component in the diagram as $E$ is exactly the excess of $E$, $e_{E}$. We have to count the dot in the same connected component as the essential set $[i,j]$ as many times as the number of inversions it is contained in, i.e., as many times as the number of $h$ such that $(i,j-i+1)\in T_{(h,\pi(h)-h+1}$. Consider the antidiagonal through the point $(i,j-i+1)$. On the right side of such a diagonal there are $r$ dots, equal to the rank of $[1,n]$. Among these points, those that satisfy the property described above are the dots not in $P_{(i,j-i+1)}$. It follows that, if the essential set $E$ has rank $r$, there are $k-r$ such points. Hence we can write:
$\ell(\pi)$ $=\sum_{(i,j)\in D(\pi)}|D(\pi)\cap T_{(i,j)}|=\sum_{i\in[n]}\sum_{\mathcal{E}\ni E\subseteq[i,\pi(i)]}e_{E}=\sum_{(r,E)\in\mathcal{E}^{r}}e_{E}(k-r).$
The last equality in the statement holds because $e_{E}>0$ if and only if $(r,E)\in\mathcal{C}^{r}$, i.e., it is in the core of the ranked essential family. ∎
@@ 5.2 Boundaries of a positroid cell
Having an explicit formula for the dimension, we can study how to construct positroid cells lying in the boundary of a given positroid cell via essential sets. In particular, we can construct all of the codimension $1$ positroid cells.
@@
Consider the positroid $\mathcal{P}$ from Example 3.3, whose ranked essential family is $\mathcal{E}^{r}=\{(1,[5,6]),(2,[1,4]),(2,[4,7]),(3,[1,8])\}$, and whose positroid cell has codimension $5$. Then there are $9$ positroid cells of codimension $1$ in $\mathcal{S}_{\mathcal{P}}$ whose ranked essential families are described in the following table. Each row contains respectively the rank, the cyclic interval, and the excess of the essential set.
@@ 5.3 Positroid varieties
Given a rank $k$ positroid on $[n]$, labeled by the bounded affine permutation $\pi$, the positroid variety $\Pi_{\pi}$ is the Zariski closure of the positroid cell $\mathcal{S}_{\pi}$ inside $\operatorname{Gr}(k,n)$. The following result generalizes [8, Proposition 2.3], which states an equivalent result for interval positroid varieties, i.e., positroids defined by rank conditions on intervals.
@@
The positroid variety $\Pi_{\pi}$ is defined as a scheme by the rank conditions $\operatorname{rk}(I)\leq r$ for every $(r,I)\in\mathcal{C}\mathcal{E}_{\pi}^{r}$, the connected elements of the ranked essential family of $\mathcal{P}_{\pi}$.
@@
By Remark 3.11 and Theorem 3.9, the rank of each cyclic interval $[i,j]$ can be computed in terms of the rank of the connected essential sets. Hence, the rank condition for $[i,j]$ is implied by those in $\mathcal{C}\mathcal{E}^{r}$. ∎
@@ 6 Small rank positroids
$\mathcal{F}^{r}=\{(r,F)\mid F\in\mathcal{F},r=\operatorname{rk}(F)<|F|\}.$
@@
If $\mathcal{P}$ is a loopless positroid of rank $2$, its ranked essential family $\mathcal{E}^{r}$ coincides with $\mathcal{F}^{r}$.
@@
If $\mathcal{P}$ has rank $2$, then $\mathcal{F}^{r}$ is the family of connected components of the graph associated to the positroid $\mathcal{P}$ as constructed in [12, Section 3.2]. In [12, Proposition 3.9], it is shown that every element of $\mathcal{F}^{r}$ is a cyclic interval, making it an essential set. Conversely, each essential set of rank $1$ gives rise to a connected component in the graph associated with the positroid $\mathcal{P}$. Hence, $\mathcal{E}^{r}=\mathcal{F}^{r}$. ∎
@@
If $\mathcal{M}$ is a loopless matroid of rank $2$, $\mathcal{M}$ is a positroid if and only if $\mathcal{F}^{r}$ consists of cyclic intervals.
@@
$\mathcal{E}^{r}=\{(2,[1,4]),(2,[4,1]),(3,[1,6])\}.$
Then $\{1,4\}$ is a circuit of rank $1$ which does not appear in $\mathcal{E}^{r}$.

### v3

@@ 5 Realization spaces of positroids
We now turn our attention to the realization space of a positroid in the totally non-negative Grassmannian and the Grassmannian. Recall that, given a field $\mathbb{K}$, the Grassmannian $\operatorname{Gr}(k,n)$ is the space of all $k$-dimensional linear subspaces of $\mathbb{K}^{n}$. Any point $V$ in $\operatorname{Gr}(k,n)$ can be represented by a $k\times n$ matrix with entries in $\mathbb{K}$. Let $M=(m_{ij})$ be a $k\times n$ matrix of indeterminates. For a subset $I=\{i_{1},\ldots,i_{k}\}\in\binom{[n]}{k}$, let $M_{I}$ denote the $k\times k$ submatrix of $M$ with the column indices $i_{1},\ldots,i_{k}$. The Plücker coordinates of $V$ are $p_{I}(V)=\text{det}(M_{I})$ for each $I\in\binom{[n]}{k}$. The Plücker coordinates do not depend on the choice of matrix $M$, up to simultaneous rescaling by a non-zero constant, and they determine the Plücker embedding of $\operatorname{Gr}(k,n)$ into $\mathbb{P}^{\binom{n}{k}-1}$. Moreover, any point $V\in\operatorname{Gr}(k,n)$ can be represented as $\text{span}\{v_{1},\ldots,v_{k}\}$ for some $\mathbb{K}$-vector space basis $\{v_{1},\dots,v_{k}\}$.
@@ 5.1 Positroid cells
Let $\mathbb{K}=\mathbb{R}$. The totally non-negative Grassmannian $\operatorname{Gr}_{\geq 0}(k,n)$ is the subset of $\operatorname{Gr}(k,n)$ consisting of points whose non-zero Plücker coordinates have all the same sign. Fix a positroid $\mathcal{P}$ on $[n]$ of rank $k$, defined via its basis $\mathcal{B}$. Then the positroid cell of $\mathcal{P}$ is the set:
$\mathcal{S}_{\mathcal{P}}=\{M\in\operatorname{Gr}_{\geq 0}(k,n)\mid M_{I}=0\text{ if and only if }I\not\in\mathcal{B}\}.$
@@
Let $\mathcal{E}^{\mathrm{r}}$ be the set of ranked essential sets of a rank $k$ positroid $\mathcal{P}$. Then
$\operatorname{codim}(\mathcal{S}_{\mathcal{P}})=\sum_{(r,[i,j])\in\mathcal{E}^{\mathrm{r}}}(k-r)e_{[i,j]}=\sum_{(r,[i,j])\in\mathcal{C}^{\mathrm{r}}}(k-r)e_{[i,j]}.$
@@
Let $\pi$ be the bounded affine permutation corresponding to the positroid $\mathcal{P}$. By Theorem 5.1, we need to prove that $\ell(\pi)=\sum_{(r,[i,j])\in\mathcal{E}^{\mathrm{r}}}e_{[i,j]}(k-r)$. Consider the dotted array of $\pi$ and note that the length of $\pi$ can be computed by looking at the array as:
@@
Note that, if the dot in row $i$ lies in $D(\pi)\cap T_{(j,\pi(j)-j+1)}$, then $(j,\pi(j)-j+1)$ defines a bounded section of the diagram of $\pi$, which is non-empty since it contains the dot $(i,\pi(i)-i+1)$. Hence there is a corner, i.e. an essential set inside $P_{(j,\pi(j)-j+1)}$. Let $(r,I)\in\mathcal{E}^{\mathrm{r}}$ be such an essential set. The number of dots in the same connected component in the diagram as $I$ is exactly the excess of $I$, $e_{I}$. We have to count the dot in the same connected component as the essential set $[i,j]$ as many times as the number of inversions it is contained in, i.e. as many times as the number of $h$ such that $(i,j-i+1)\in T_{(h,\pi(h)-h+1}$. Consider the antidiagonal through the point $(i,j-i+1)$. On the right side of such a diagonal there are $r$ dots, equal to the rank of $[1,n]$. Among these points, those that satisfy the property described above are the dots not in $P_{(i,j-i+1)}$. It follows that, if the essential set $I$ has rank $r$, there are $k-r$ such points. Hence we can write:
$\ell(\pi)$ $=\sum_{(i,j)\in D(\pi)}|D(\pi)\cap T_{(i,j)}|=\sum_{i\in[n]}\sum_{\mathcal{E}\ni I\subseteq[i,\pi(i)]}e_{I}=\sum_{(r,I)\in\mathcal{E}^{\mathrm{r}}}(k-r)e_{I}.$
The last equality in the statement holds because $e_{I}>0$ if and only if $(r,I)\in\mathcal{C}^{\mathrm{r}}$, i.e. it is in the core of the ranked essential family. ∎
@@ 5.2 Boundaries of a positroid cell
Having an explicit formula for the dimension, we can study how to construct positroid cells lying in the boundary of the Zariski closure of a given positroid cell via essential sets. In particular, we can construct all of the codimension $1$ positroid cells.
@@
Consider the positroid $\mathcal{P}$ from Example 3.3, whose ranked essential family is $\mathcal{E}^{\mathrm{r}}=\{(1,[5,6]),(2,[1,4]),(2,[4,7]),(3,[1,8])\}$, and whose positroid cell has codimension $5$. Then there are $9$ positroid cells of codimension $1$ contained in the Zariski closure of $\mathcal{S}_{\mathcal{P}}$ whose ranked essential families are described in the following table. Each row contains respectively the rank, the cyclic interval, and the excess of the essential set.
Figure 4: The positroids whose cell lies in the Zariski closure of the cell of the positroid from Example 5.4. The data on the left of each box is the rank, the cyclic interval and the excess of each essential set of the family. On the right, the point-line configuration of the positroid is depicted.
@@ 5.3 Positroid varieties
Given a rank $k$ positroid on $[n]$, labeled by the bounded affine permutation $\pi$, the positroid variety $\Pi_{\pi}$ is the Zariski closure of the positroid cell $\mathcal{S}_{\pi}$ inside $\operatorname{Gr}(k,n)$. The following result generalizes [9, Proposition 2.3], which states an equivalent result for interval positroid varieties, i.e. positroids defined by rank conditions on intervals.
@@
The positroid variety $\Pi_{\pi}$ is defined as a scheme by the rank conditions $\operatorname{rk}(I)\leq r$ for every $(r,I)\in\mathcal{C}\mathcal{E}_{\pi}^{\mathrm{r}}$, the connected elements of the ranked essential family of $\mathcal{P}_{\pi}$.
@@
By Remark 3.11 and Theorem 3.9, the rank of each cyclic interval $[i,j]$ can be computed in terms of the rank of the connected essential sets. Hence, the rank condition for $[i,j]$ is implied by those in $\mathcal{C}\mathcal{E}^{\mathrm{r}}$. ∎
@@ 6 Small rank positroids
$\mathcal{F}^{\mathrm{r}}=\{(r,F)\mid F\in\mathcal{F},r=\operatorname{rk}(F)<|F|\}.$
@@
If $\mathcal{P}$ is a loopless positroid of rank $2$, its ranked essential family $\mathcal{E}^{\mathrm{r}}$ coincides with $\mathcal{F}^{\mathrm{r}}$.
@@
If $\mathcal{P}$ has rank $2$, then $\mathcal{F}^{\mathrm{r}}$ is the family of connected components of the graph associated to the positroid $\mathcal{P}$ as constructed in [13, Section 3.2]. In [13, Proposition 3.9], it is shown that every element of $\mathcal{F}^{\mathrm{r}}$ is a cyclic interval, making it an essential set. Conversely, each essential set of rank $1$ gives rise to a connected component in the graph associated with the positroid $\mathcal{P}$. Hence, $\mathcal{E}^{\mathrm{r}}=\mathcal{F}^{\mathrm{r}}$. ∎
@@
If $\mathcal{M}$ is a loopless matroid of rank $2$, $\mathcal{M}$ is a positroid if and only if $\mathcal{F}^{\mathrm{r}}$ consists of cyclic intervals.
@@
$\mathcal{E}^{\mathrm{r}}=\{{(1,[1,2]),(2,[1,5]),(2,[5,2]),(3,[1,7])}\}.$
Then $\{1,5\}$ is a circuit of rank $1$ which does not appear in $\mathcal{E}^{\mathrm{r}}$. Moreover, $\{1,2\}$ is not a flat of the positroid, as its closure is given by $\{1,2,5\}$.